OpenStudy (anonymous):

A box of ground nutmeg weighs 1 1/3 ounce. If there are 20 teaspoons of nutmeg in the box, how much does one teaspoon of nutmeg weigh? 1) 1/15 ounce 2) 1/20 ounce 3) 1/60 ounce 4) 4/3 ounce

OpenStudy (anonymous):

i guess the box weighs nothing?

OpenStudy (anonymous):

i think this is a lame attempt at a word problem asking you to divide \(1\tfrac{1}{3}\) by \(20\)

OpenStudy (anonymous):

i.e. compute \[1\tfrac{1}{3}\div 20=\frac{4}{3}\div 20=\frac{4}{3\times 20}=\frac{1}{3\times 5}\]

OpenStudy (anonymous):

This is not helping me...

OpenStudy (anonymous):

your job is to divide \[1\tfrac{1}{3}\] by \(20\) do you know how to do it?

OpenStudy (anonymous):

No

OpenStudy (anonymous):

I Dont Have A calculator at hand atm.

OpenStudy (anonymous):

first step it to change the mixed number \(1\tfrac{1}{3}\) to an improper fraction do you know how to do that?

OpenStudy (anonymous):

i think the point is to learn how to do this without a calculator of course since you are on a computer you have an unlimited amount of on line calculators you can use, but since you are working with fractions the idea is to learn how to divide a mixed number by a whole number

OpenStudy (anonymous):

can you turn \[1\tfrac{1}{3}\] into an improper fraction?

OpenStudy (anonymous):

How did you come up with 1 over 3x5 satellite73?